If f(X,y)={xy/x2+y2, if(X,y)#(0,0)

0 if (X,y)=(0,0)
Check the continuity CF F(X,y) at (0,0)
1
Expert's answer
2020-04-27T17:43:45-0400

I have doubts in the question, so I have given two answers.

If the given function is

"f(x)=\begin{cases}
\frac{xy}{2x+2y}&\text{if} (x,y)\neq(0,0) \\
0 & \text{if} (x,y)=(0,0)
\end{cases}"


Consider the sequence "x_n=(\frac{1}{n},\frac{-1}{n})" in "\R^2" .

Then "\text{lim}_{n\to \infin}x_n=(0,0)." But "f(x_n)=\frac{( \frac{1}{n})( \frac{-1}{n} )}{\frac{2}{n}-\frac{2}{n}}"

"\implies f(x_n)=\frac{-1}{0}" .

which does not exist.

Hence "f(x,y)" is not continuous at (0,0).


If the function is "f(x)=\begin{cases}
\frac{xy}{x^2+y^2} & \text{if} (x,y)\neq(0,0) \\
0 & \text{if} (x,y)=(0,0)
\end{cases}"


Consider the sequence "x_n=(\frac{1}{n},\frac{1}{n}) \ \text{in} \ \R^2."

Then "\text{lim}_{x\to \infin} x_n=(0,0)." But "f(x_n)=\frac{ \frac{1}{n^2}}{\frac{2}{n^2}}=\frac{1}{2}" ,which is a constant function and converge to "\frac{1}{2}" .

"\therefore f(x_n)\nrightarrow f(0)" as "n\rightarrow \infin" .

Hence "f(x,y)" is not continuos at (0,0).